# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ III An algorithm for converting QSP parameters to AIM
Here, we present a method for aligning the phases of the QSP sequence with those in AIM. We develop an algorithm to adapt QSP phases for AIM, which provides the experimental inputs—amplitude, frequency, and timings—needed for the direct realization of the QSP sequence on an experimental device with LZSM-type evolution, using the method of determining parameters from AIM [14].
We now consider a harmonic signal that is linear in the vicinity of the anti-crossing region, meaning it corresponds to analytical solutions for the probability of an LZSM transition and results in the LZSM transition probability. From Eq. (14c) we have
@@
The total phase gain during the LZSM transition is called the Stückelberg phase and it is expressed as the sum of the Stokes phase $\phi_{\text{S}}$ given by the Eq. (14b), gained during diabatic evolution, and the phase accumulated during the adiabatic evolution of the system $\zeta(t_{i},t_{j})$ given by Eq. (15a)
[…]
The steps for determining the amplitude $A$, frequency $\omega$, and time $\Delta t=t_{j}-t_{i}$ for a driven signal with a LZSM transition can be described as follows:
1. Given that the system undergoes a rotation by an angle $\theta$ about the $x$-axis, the probability of finding the system in the upper state after a single LZSM transition can be found from the following equation:
@@
Next, since we start and finish the transition far from the anti-crossing point, we fix the amplitude to ensure that the next transition starts from the same detuning $\varepsilon$ as when we finish the previous one. Using the previously defined LZSM probability in Eq. (14c), we obtain the driving frequency:
@@
2. The phase correspondence condition can be found by comparing the evolution matrices for a two-level system described by AIM in Eqs. (14a, 14b, 16) and the one for QSP with corresponding QSP phases in Eq. (4):
@@
3. We use a signal, shown in Fig. 2 and defined as,
@@
Figure 2: Signal shape for the BB1 sequence realized with AIM. We assumed large amplitude, $A=3\Delta$. The blue line corresponds to the harmonic driven evolution described by the cosine in Eq. (23), the orange line corresponds to the adjustable phase gain while the driving is constant, which is used to match the QSP phases with the AIM phases, and the red crosses are the transition points. Time is normalized to the qubit resonant excitation period $T_{\text{r}}$.
@@
... (24)
$2\phi_{\text{S}}+2\zeta_{\text{const},d-1}(t)+2\zeta_{\text{cos}}(t)+\pi+2\pi j=-2\phi_{d-1},$
$\phi_{\text{S}}+2\zeta_{\text{const},d}(t)+2\zeta_{\text{cos}}(t)+\pi+2\pi j=-2\phi_{d},\hskip 1.99997pt\text{$d$ is even,}$,
$\phi_{\text{S}}+2\zeta_{\text{const},d}(t)+2\zeta_{\text{cos}}(t)+2\pi j=-2\phi_{d},\hskip 10.00002pt\text{$d$ is odd,}$,
here $j$ is an integer and $k=0,1,...,d$,
$\zeta_{\text{const},k}=\Omega_{\text{L}}T_{\text{const},k}/2,\hskip 10.00002pt\Omega_{\text{L}}=\sqrt{A^{2}+\Delta^{2}}/\hbar,$ (25a)
[…]
@@
where $T_{\text{const},k}$ is the duration of the free evolution stage, $T_{\text{cos}}=2\pi/\omega$ is the period of the resonance driving. Note that free evolution precesion around the $z$-axis happen with the Larmor frequency $\Omega_{\text{L}}$ and period $T_{\text{L}}=2\pi/\Omega_{\text{L}}$. In these conditions, the phases $\zeta_{\text{const}}$ and $\zeta_{\text{cos}}$ are gained during the free-evolution stage, where $\varepsilon=\pm A$, and during the LZSM transition, where $\varepsilon=\pm A\text{cos}(\omega t)$, respectively. After determining the target phases $\boldsymbol{\Phi}=(\phi_{0},\phi_{1},\dots,\phi_{d})$, the evolution time is obtained to satisfy the conditions in Eq. (III) and achieve the target QSP parameters mentioned above.
@@
In this section we demonstrate the direct implementation of the QSP algorithm evolution using AIM. The execution time of the operations is determined by the LZSM probability in Eq. (14c), which corresponds to the rotation angle around the $x$-axis on the Bloch sphere and the chosen amplitude of the driving signal $A$. Notably, for specific values of the LZSM probability $\mathcal{P}=0$, the operation time $T_{\text{cos}}=2\pi/\omega$, see Eq. (21). diverges to infinity. The next section is devoted to a way how to resolve this problem.
IV Double Landau-Zener-Stückelberg-Majorana transition quantum signal processing algorithm
@@
Here we describe how to use double LZSM transitions [14] for realizing QSP. Double LZSM transitions allow eliminating the huge time variability of each QSP rotation around the $x$-axis. A double LZSM transition is fully equivalent to a Mach-Zehnder interferometer with two beam splitters (LZSM transitions) with a phase gain between beam splitters or LZSM transitions and interferometry between states [36, 13, 37] (adiabatic evolution). Here we use $\mathcal{P}=0.5$, which is equivalent to a (commonly used in photonics) 50/50 beam splitter.
@@
In particular, the Mach-Zehnder interferometer in photonic quantum computers is used as a programmable quantum gate, programmed by adjusting the phase gain between beam splitters [38, 39, 40]. To adjust the rotation angle of the signal rotation operator $W$, we use adjustable additional free evolution between two LZSM transitions. For a double LZSM transition [14] with adjustable phase gain between two transitions, we have
$\Xi$ $=$ $U_{1}N_{2}U_{1}U_{2}N_{1}^{\text{inv}}U_{2}$ (26)
$=$ $U^{\text{inv}}_{\text{LZSM(1)}}U_{\text{LZSM(2)}}=\begin{pmatrix}\Xi_{11}&\Xi_{12}\\ -\Xi_{12}^{*}&\Xi_{11}^{*}\end{pmatrix},$
@@
where for the adiabatic evolution between transitions we use $U_{2}$ with an adjustable phase, $U_{2}=U(\zeta_{2}+\phi_{\text{ad}})$, $\phi_{\text{ad}}$ is an additional phase gain between two LZSM transitions. For the Stückelberg phase $\phi_{\text{St}}=\phi_{\text{S}}+\zeta_{2}$, we choose conditions for the X gate from Ref. [14], $\phi_{\text{St}}=\frac{\pi}{2}$, $\mathcal{R}=\mathcal{T}={\frac{1}{\sqrt{2}}}$. Then we obtain
$\Xi_{11}$ $=[\mathcal{R}^{2}\exp{(-i(2\phi_{\text{S}}+2\zeta_{2}+2\phi_{\text{ad}}))}+\mathcal{T}^{2}]$
@@
$\Xi_{12}$ $=\mathcal{R}\mathcal{T}[1-\exp{(-i(2\phi_{\text{S}}+2\zeta_{2}+2\phi_{\text{ad}}))}]$
@@
$\zeta_{2}$ $\!\!\!\!+\phi_{\text{ad}}=\zeta(t_{S1},t_{S2}),$
@@
Next, to obtain the parameters for rotation around the x-axis on a Bloch sphere, we compare the double LZSM transition matrix elements with the respective ones for a rotation operator around the $x$-axis on a Bloch sphere $R_{x}(\theta)$
@@
Then to obtain the phases, necessary to perform the $R_{x}$ gate, we solve the corresponding system of equations $\Xi=R_{x}(\theta)$:
$\begin{cases}\cos\frac{\theta}{2}=\frac{1}{2}\left(1+e^{-i(\pi+2\phi_{\text{ad}})}\right)e^{i(\phi_{\text{ad}}-\phi_{\text{in}}-\phi_{\text{fin}})}\\ e^{i3\pi/2}\sin\frac{\theta}{2}=\frac{1}{2}\left(1-e^{-i(\pi+2\phi_{\text{ad}})}\right)e^{i(\phi_{\text{ad}}+\phi_{\text{in}}-\phi_{\text{fin}}+\pi/2)}\end{cases}.$ (31)
As a result, by using Euler’s formula for cosine and sine, we obtain a system of equations for the phases. By solving this system of equations, we obtain additional phases to perform the $R_{x}(\theta)$ operation for the same LZSM driving parameters as for the $X$ gate in Ref. [14]
@@
$\begin{cases}t_{\text{ad}}=\left\{\dfrac{\theta+\pi}{2\pi}\right\}T_{\text{L}},\\ t_{\text{in}}=\left\{\dfrac{\pi/2}{2\pi}\right\}T_{\text{L}}=T_{\text{L}}/4,\\ t_{\text{fin}}=\left\{\dfrac{5\pi/2}{2\pi}\right\}T_{\text{L}}=T_{\text{L}}/4.\end{cases}$ (33)
@@
With the driving signal $\varepsilon(t)=-A\cos\omega t$, the first transition will be reversed, that lead to an additional $\pi$ phase to the initial phase gain [14]. This gives the free evolution time that should be added at the beginning of the sequence, denoted as $t_{\text{in}}$, and between the signal rotation operators, denoted as $t_{\text{mid}}$ for the rotation around the $x$-axis
$t_{\text{mid}}=\left\{\frac{5\pi/2+\pi/2}{2\pi}\right\}T_{\text{L}}=T_{\text{L}}/2.$ (34)
@@
Figure 4: Panel (a) shows the resulting occupation probability $P_{-}$ for the BB1 sequence in Eq. (12) compared to the qubit dynamics simulation result with the driving signal defined by AIM. Panel (b) presents a comparison of the hardware execution time for the BB1 sequence using different approaches: (1) IBM-Q time (blue line), computed as the difference between the execution time with zero gates and the time with the BB1 sequence. This isolates the time required solely for the QSP algorithm, enabling direct comparison with other methods. (2) IBM-Q with the highest optimization setting (orange line), where the execution time does not depend on the number of single-qubit gates—i.e., the entire BB1 sequence is replaced with a single compiled gate. (3) Direct realization of QSP via AIM (green curve). (4) Double LZSM transitions (purple line). To compare times we divide all of them by the period of the resonance driving $T_{\text{r}}={2\pi}/{\Delta}$.
@@
So, we described how to get all the necessary phases such as the phase gain time-intervals before the sequence, between double LZSM transitions, and after the sequence. To get the amplitude and the frequency we use the algorithm, described in Ref. [14] for the X-gate. With these we have all the parameters required to implement QSP sequences using double LZSM transitions.
@@
Double LZSM transitions allow us to eliminate the main weakness of the direct implementation of the QSP algorithm with AIM: the transition time diverges as the signal rotation angle approaches zero, making the result highly susceptible to decoherence and dissipation. Double transitions allow to perform the QSP sequence fast and with high fidelity by leveraging the LZSM gates technique, using the same driving signal defined for the X-gate in Ref. [14] with the additional free evolution time between transitions $t_{\text{ad}}$ for the signal operator $W=R_{x}(\theta)$. Then we perform the signal processing operator $S=R_{z}(\phi_{i})$ after each transition. By choosing a working point with higher amplitude, we can significantly improve the fidelity of LZSM gates by slightly decreasing the gate execution time.
@@
Figure 3(a) shows the driving signal used to implement the $R_{x}(\theta)$ gate, where the time between transitions $t_{\text{ad}}$ (upper panel) is defined by the rotation angle around the $x$-axis, and the time intervals $t_{\text{in}}$ and $t_{\text{fin}}$ are used to gain additional phases before and after transitions to perfectly match our double LZSM transition matrix to an $R_{x}(\theta)$ gate.
Figure 3(b) shows the equivalent Mach-Zehnder interferometer scheme which is used as a programmable quantum logic gate for photons, where the rotation angle around the $x$-axis is defined by additional phase gain in one of the optical ways of the Mach-Zehnder interferometer, that can be tunable.

### v3

@@ III An algorithm for converting QSP parameters to AIM
Here, we present a method for aligning the phases of the QSP sequence with those in AIM. We develop an algorithm to adapt QSP phases for AIM, which provides the experimental inputs—amplitude, frequency, and timings—needed for the direct realization of the QSP sequence on an experimental device with LZSM-type evolution [20], using the method of determining driving parameters for AIM [18].
The harmonic signal $\varepsilon(t)=-A\cos\omega t$ is linear in the vicinity of the anti-crossing region, meaning it corresponds to analytical solutions for the probability of an LZSM transition and results in the LZSM transition probability. From Eq. (14f) we have
@@
The total phase gain during the LZSM transition is called the Stückelberg phase and it is expressed as the sum of the Stokes phase $\phi_{\text{S}}$ given by Eq. (14e), gained during diabatic evolution, and the phase accumulated during the adiabatic evolution of the system $\zeta(t_{i},t_{j})$ given by Eq. (15a)
$\phi_{\text{St}}=\phi_{\text{S}}+\zeta(t_{i},t_{j}).\hskip 10.00002pt$ (19)
The steps for determining the amplitude $A$, frequency $\omega$, and time $\Delta t=t_{j}-t_{i}$ for a driven signal with an LZSM transition can be described as follows:
1. Given that the system undergoes a rotation by an angle $\theta$ about the $x$-axis, the probability of finding the system in the upper state after a single LZSM transition can be found from the following expression:
@@
Next, since we start and finish the transition far from the anti-crossing point, we fix the amplitude to ensure that the next transition starts from the same detuning $\varepsilon$ as when we finish the previous one. Using the previously defined LZSM probability in Eq. (14f), we obtain the driving frequency:
@@
2. The phase correspondence condition can be found by comparing the evolution matrices for a two-level system described by AIM in Eqs. (14a, 14c, 16) and the one for QSP with corresponding phases in Eq. (4):
@@
3. We use a signal, shown in Fig. 2 and defined as
@@
Figure 2: Signal shape for the BB1 sequence realized with AIM and a large amplitude $A>\Delta$. The blue line corresponds to the harmonic-driven evolution described by the cosine in Eq. (23), the orange line corresponds to the adjustable phase gain while the driving is constant, which is used to match the QSP phases with the AIM phases, and the red crosses are the transition points. Time is normalized to the qubit resonant excitation period $T_{\text{q}}={2\pi}/{\Delta}$.
@@
$2\phi_{\text{S}}+2\zeta_{\text{const},k}(t)+2\zeta_{\text{cos}}(t)+\pi+2\pi j=-2\phi_{k},$ (24)
$\phi_{\text{S}}+2\zeta_{\text{const},d}(t)+2\zeta_{\text{cos}}(t)+\pi+2\pi j=-2\phi_{d},\hskip 1.99997ptd\text{ is even,}$
$\phi_{\text{S}}+2\zeta_{\text{const,}d}(t)+2\zeta_{\text{cos}}(t)+2\pi j=-2\phi_{d},\hskip 10.00002ptd\text{ is odd.}$
Here $j$ is an integer and k=0,1,,dk=0,1,...,d,
$\zeta_{\text{const},k}=\Omega_{\text{L}}T_{\text{const},k}/2,$ (25a)
$\Omega_{\text{L}}=\sqrt{A^{2}+\Delta^{2}}/\hbar,$ (25b)
$\zeta_{\text{cos}}=\dfrac{1}{2\hbar}\int_{0}^{T_{\text{cos}}/2}\sqrt{A^{2}\cos^{2}\big(\omega(t-T_{\text{s}})\big)+\Delta^{2}}dt,$ (25c)
where $T_{\text{const},k}$ is the duration of the free evolution stage, $T_{\text{cos}}=2\pi/\omega$ is the period of the resonance driving.
@@
Note that free evolution precession around the $z$-axis happens with the Larmor frequency $\Omega_{\text{L}}$ and period $T_{\text{L}}=2\pi/\Omega_{\text{L}}$. In these conditions, the phases $\zeta_{\text{const}}$ and $\zeta_{\text{cos}}$ are gained during the free-evolution stage, where $\varepsilon=\pm A$, and during the LZSM transition, where $\varepsilon=\pm A\text{cos}(\omega t)$, respectively. After determining the target phases $\boldsymbol{\Phi}~=~(\phi_{0},\phi_{1},\dots,\phi_{d})$, the evolution time is obtained to satisfy the conditions in Eq. (III) and achieve the target QSP parameters mentioned above.
@@
To summarize, in this section, we demonstrated the direct analogy of the QSP technique evolution using AIM. The execution time of the operations is determined by the LZSM probability in Eq. (14f), which corresponds to the rotation angle around the $x$-axis on the Bloch sphere and the chosen amplitude of the driving signal $A$. Notably, for specific values of the LZSM probability $\mathcal{P}=0$, the operation time $T_{\text{cos}}=2\pi/\omega$, see Eq. (21) diverges to infinity. The next section is devoted to a way to resolve this problem.
IV quantum signal processing
based on double Landau-Zener-Stückelberg-Majorana transitions
@@
Here we describe how to use double LZSM transitions for realizing QSP. This allows eliminating the huge time variability of each QSP rotation around the $x$-axis. A double LZSM transition is fully equivalent to a Mach-Zehnder interferometer with two beam splitters (LZSM transitions) with a phase gain between beam splitters or LZSM transitions and interferometry between states (adiabatic evolution) [42, 17, 43]. Here we use $\mathcal{P}=0.5$, which is equivalent to a (commonly used in photonics) 50/50 beam splitter.
@@
In particular, the Mach-Zehnder interferometer in photonic quantum computers is used as a controllable quantum gate, controlled by adjusting the phase gain between beam splitters [44, 45, 46]. To adjust the rotation angle of the signal operator $W$, we use an adjustable additional free evolution between two LZSM transitions. For a double LZSM transition [17] with adjustable phase gain between two transitions, we have
$\Xi$ $=$ $U_{2}N_{2}U_{2}U_{1}N_{1}^{\text{inv}}U_{1}$ (26)
$=$ $U_{\text{LZSM(2)}}U^{\text{inv}}_{\text{LZSM(1)}}=\begin{pmatrix}\Xi_{11}&\Xi_{12}\\ -\Xi_{12}^{*}&\Xi_{11}^{*}\end{pmatrix}$,
@@
where for the adiabatic evolution between transitions we use $U_{2}$ with an adjustable phase, $U_{2}=U(\zeta_{2}+\phi_{\text{ad}})$, where $\phi_{\text{ad}}$ is an additional phase gain between two LZSM transitions. For the Stückelberg phase $\phi_{\text{St}}=\phi_{\text{S}}+\zeta_{2}$, we choose conditions for the X gate from Ref. [18], $\phi_{\text{St}}=\frac{\pi}{2}$, $\mathcal{R}=\mathcal{T}={\frac{1}{\sqrt{2}}}$. Then we obtain
$\Xi_{11}=[\mathcal{R}^{2}\exp{(-i(2\phi_{\text{S}}+2\zeta_{2}+2\phi_{\text{ad}}))}+\mathcal{T}^{2}]$ (27a)
@@
$\Xi_{12}=\mathcal{R}\mathcal{T}[1-\exp{(-i(2\phi_{\text{S}}+2\zeta_{2}+2\phi_{\text{ad}}))}]$ (27b)
@@
$\zeta_{2}+\phi_{\text{ad}}=\zeta(t_{S1},t_{S2}),$ (28)
@@
Next, to obtain the parameters for rotation around the x-axis on a Bloch sphere, we compare the double LZSM transition matrix elements with the respective ones for a rotation operator $R_{x}(\theta)$ on a Bloch sphere
@@
Then to obtain the phases, needed to perform the $R_{x}(\theta)$ gate, we use the same driving parameters as in the X gate in Ref. [18] and obtain the corresponding system of equations $\Xi=R_{x}(\theta)$:
$\begin{cases}\cos\frac{\theta}{2}=\frac{1}{2}\left(1+e^{-i(\pi+2\phi_{\text{ad}})}\right)e^{i(\phi_{\text{ad}}-\phi_{\text{in}}-\phi_{\text{fin}})},\\[8.61108pt] e^{i3\pi/2}\sin\frac{\theta}{2}=\frac{1}{2}\left(1-e^{-i(\pi+2\phi_{\text{ad}})}\right)e^{i(\phi_{\text{ad}}+\phi_{\text{in}}-\phi_{\text{fin}}+\pi/2)}.\end{cases}$ (30)
As a result, by using Euler’s formula, we obtain a system of equations for the phases. By solving this system of equations, we obtain additional phases to perform the $R_{x}(\theta)$ operation
@@
$\begin{cases}t_{\text{ad}}=\left\{\dfrac{\theta+\pi}{2\pi}\right\}T_{\text{L}},\\ t_{\text{in}}=\left\{\dfrac{\pi/2}{2\pi}\right\}T_{\text{L}}=\dfrac{T_{\text{L}}}{4},\\ t_{\text{fin}}=\left\{\dfrac{5\pi/2}{2\pi}\right\}T_{\text{L}}=\dfrac{T_{\text{L}}}{4}.\end{cases}$ (32)
@@
With the driving signal $\varepsilon(t)=-A\cos\omega t$, the first transition will be reversed, and leads to an additional $\pi$ phase to the initial phase gain [17]. This gives the free evolution time that should be added at the beginning of the sequence, denoted as $t_{\text{in}}$, and between the signal operators, denoted as $t_{\text{mid}}$ for the rotation around the $x$-axis
$t_{\text{mid}}=\left\{\frac{5\pi/2+\pi/2}{2\pi}\right\}T_{\text{L}}=\frac{T_{\text{L}}}{2}.$ (33)
@@
Figure 4: Panel (a) shows the resulting occupation probability $P_{-}$ for the BB1 sequence in Eq. (12) compared to the qubit dynamics simulation result with the driving signal defined by AIM. Panel (b) presents a comparison of the hardware execution time $t/T_{\text{q}}$ for the BB1 sequence using different approaches: (1) IBM-Q time (blue line), computed as the difference between the execution time with zero gates and the time with the BB1 sequence. This isolates the time required solely for the QSP technique, enabling direct comparison with other methods. (2) IBM-Q with the optimization (orange line), where the execution time does not depend on the number of single-qubit gates—i.e., the entire BB1 sequence is replaced with a single compiled gate. (3) Direct realization of QSP via AIM (green curve). (4) Double LZSM transitions (purple line). To compare execution times we divide all of them by the corresponding period of the qubit resonance driving $T_{\text{q}}={2\pi}/{\Delta}$.
@@
So, we described how to get all the necessary phases, such as the phase gain time-intervals before the sequence, between double LZSM transitions, and after the sequence. To obtain the amplitude and the frequency, we use the algorithm described in Ref. [18] for the X gate. With these, we have all the parameters required to implement QSP sequences using double LZSM transitions.
@@
Double LZSM transitions allow us to eliminate the main weakness of the direct implementation of the QSP technique with AIM: the transition time diverges as the signal rotation angle approaches zero, making the result highly susceptible to decoherence and dissipation. Double transitions allow to perform the QSP sequence fast and with high fidelity by leveraging the LZSM gates technique, using the same driving signal defined for the X gate with the additional free evolution time between transitions $t_{\text{ad}}$ for the signal operator $W=R_{x}(\theta)$. Then we apply the signal processing operator $S=R_{z}(\phi_{i})$ after each transition. By choosing a working point with a higher amplitude, we can significantly improve the fidelity of LZSM gates by slightly decreasing the gate execution time.
@@
Figure 3(a) shows the driving signal used to implement the $R_{x}(\theta)$ gate, where the time between transitions $t_{\text{ad}}$ is defined by the rotation angle around the $x$-axis, and the time intervals $t_{\text{in}}$ and $t_{\text{fin}}$ are used to gain additional phases before and after transitions to perfectly match our double LZSM transition matrix to an $R_{x}(\theta)$ gate. Figure 3(b) shows the equivalent Mach-Zehnder interferometer scheme, which is used as a controllable quantum logic gate for photons, where the rotation angle around the $x$-axis is defined by additional phase gain in one of the optical ways of the Mach-Zehnder interferometer, which can be tunable.